Time of last reading:
March 20, 2019, 3 p.m.
Dec. 20, 2017, 4:45 a.m.
Issue Feb. 12, 2018 - Feb. 22, 2018
It has been brought to our attention by Environmental Instruments Ltd (manufacturers of AQMesh) that due to a bug in a server systems update that NO2 and O3 data from all AQMesh pods were effected from 7am on 12th February until 9am on 22nd February 2018. No other channels are affected. We caution you against using NO2 and O3 data for this period until the corrections have been applied. We urge users to re-call this data from Tuesday 27th when all readings will have been re-processed. The Urban Observatory apologises for the inconvenience.Issue May 1, 2018 - May 1, 2018
Aq_meshes 1904 (Chillingham Rd Primary School) , 1905 (Corner House, Heaton) removed for slope/offset adjustment at Precision SiteIssue May 22, 2018 - May 22, 2018
Chillingham Road Primary (1904150) and Coast Road Corner House (1095150) were not 'refurbished' back in Jan 18 with the rest of that batch. These have been removed and substituted for new AQ Mesh pods ie Chillingham Road Primary is now aq_mesh1970150 Coast Road Corner House is now aq_mesh1971150 . Please refer to those pods from this date till further notice. ThanksIssue May 22, 2018 - May 22, 2018
Deployed Chillingham Road Primary School aq_mesh1970150 as replacement for aq_mesh1904150 Lamp post 30C130 Postcode NE6 5XX GPS 54.98430885 -1.576213539Issue Aug. 2, 2018 - Aug. 2, 2018
Redeployed former SenseMyStreets (was Chillingham Road Primary) Deployed at Cragside Primary School, High Heaton Lamp post 46N43 NE7 7EL 54.998579 -1.588521Issue Sept. 12, 2018 - Sept. 12, 2018
Temporary re-location for calibration at Coast Road Precision StationIssue May 4, 2018 - May 22, 2018
Moved to co-locate with Gosforth Precision cabinet During this period, no data was sent - Pod was deemed to be unreliable due to Firmware issues and a decision was made to substitue Pod aq_mesh197015 at the original Chillingham Rd primary school until these issues could be resolved by Air Monitors.Issue Oct. 10, 2018 - Oct. 10, 2018
aq_mesh1904150 collocated at Precision Station NORTH_TYNESIDE 1048100 (From 21/09/18 to 01/10/2018) CO(Linear regressions analysis can not be done due to CO analyser failure) Original slope 0.927299976 Original Offset 201.772995 New Slope New Offset R square NO Original slope 1 Original Offset -0.884599984 New Slope 0.996266049 New Offset -4.261937849 R square 0.989184702 NO2 Original slope 1 Original Offset 7.003600121 New Slope 1.129770876 New Offset 6.59404687 R square 0.827700521 O3 Original slope 0.559400022 Original Offset 9.57859993 New Slope 0.241099158 New Offset 17.04976889 R square 0.421862823 PM10 Original slope 1 Original Offset 0 New Slope 2.022510846 New Offset 4.994688706 R square 0.148116979 PM2.5 Original slope 1 Original Offset 0 New Slope 0.808569041 New Offset 2.804339121 R square 0.100319475 PM1 Original slope 1 Original Offset 0 New Slope 0.447000546 New Offset 1.843895641 R square 0.063737029Issue Oct. 24, 2018 - Oct. 25, 2018
All AQ Meshes did not report any data between Weds 24th Oct 06:00 – Thurs 25th Oct 11:15, reason not yet knownIssue Nov. 16, 2018 - Nov. 16, 2018
Approx 15:00 - 16:00 : Changed the sampling frequency for gases for all AQ Meshes to be standard across all pods : P1 (Gas Reading Frequency) - Ten secs (unchanged) P2 (Gas Averaging Frequency) - One Minute (was mostly 15 mins) P3 (Gas Transmission Frequency) - 30 Mins (was One Hour)Issue Nov. 22, 2018 - Nov. 23, 2018
22/11/18 17:00 AirMonitors carrying out server maintenance at approx 17:00 (UTC) on Thursday November 22nd 23/11/10:00 server back. Possible data loss during this periodIssue Dec. 6, 2018 - Dec. 6, 2018
06/12/2018 Major disruption to the 02 Cellphone/Data network today. This will affect many sensors that rely on that network (eg Emotes)Issue Nov. 30, 2018 - Dec. 12, 2018
Gas sensor type:O3 Sensor failed at 09.49 on 30/11/2018 Failed sensor serial number:204100706 (Date Code 117) Serial number of new sensor:204640838 (Date Code 058) Notified 05/12/18 Received 11/12/18 Replaced 12/12/18 Please allow 2-3 days for new sensor to stabiliseIssue Jan. 20, 2019 - Jan. 21, 2019
20/01/19 17:00 approx Sensors data not being displayed on the portal. This was due to failed video data stuck in the queue and blocking all other data.Issue Jan. 15, 2019 - Jan. 16, 2019
15/01/19 02:15 - 16/01/19 Data lost from all sensors due to quue failure. This data may eventually get back-filled.
Sensor Maintenance Records: